The NAV of a mutual fund is $10 per share. You invest $100,000 in the fund. The front end load is 2%. You sell your shares within six months at an NAV of $11 per share. The redemption fees is 3%. What is your return from this mutual fund investment?
Answer #1

rate positively ..

Amount invested = 100000
Front end load @ 2% = 2000
Net amount invested = 98000
NAV per share = 10
Number of share holding = 98000/10 9800
Sales price = 9800*11 107800
Redemption fees @ 3% 3234
Net sales price = 104566
Therefore return = 104566-100000 4566
Return % = 4566/100000 4.57%
Ans = 4.57%